# Signed Networks: theory, methods, and applications

### v1

@@ 8 Conclusions
Some fundamental quantities still lack accepted signed versions (e.g. assortativity). Others admit multiple possible generalisations: the Laplacian can be either opposing or repelling, node degree can be positive or negative depending on whether one considers signed, net or total degree, suitable for different contexts. In many cases, familiar properties break down when signs are introduced: shortest-path distances no longer satisfy metric axioms, and random-walk transition rates can become negative [tianSpreadingStructuralBalance2024]. Even when a measure can be generalised, its interpretation often changes fundamentally. For instance, Katz centrality describes power relations and may become negative, and subgraph centrality encodes local balance rather than importance [diaz-diazMathematicalModelingLocal2024]. Developing principled frameworks for when to use which generalisation, and clarifying their distinct interpretations, would help standardise practice across the field.
@@
The term “polarisation” is used ambiguously in signed network research, encompassing distinct phenomena. Elite polarisation (in legislative or parliamentary contexts) differs from mass polarisation (in online platforms). Structural polarisation (the emergence of differentiated groups in relationship patterns) differs from affective polarisation (emotional or attitudinal hostility between groups) [iyengar2012affect]. Moreover, signed networks enable distinguishing between mere separation, cohesive but disconnected groups, and explicit antagonism through negative ties. Understanding how these definitions relate and how empirical results differ when only endorsing actions are considered versus when negative interactions are also included is crucial [candellone2025negative], and what each reveals about underlying social or political processes remains an important open question.
@@
A foundational question remains often overlooked: what kinds of systems are best modelled as signed networks? While the framework is widely used, its application sometimes lacks clear justification. Not all oppositional or divergent relationships are best captured via edge signs. For example, there are serious concerns about treating correlation matrices as signed networks without prior processing [peixoto_NatComm_2022]. Distinguishing between situations where signed networks provide genuine insight versus where other representations (multilayer networks, attributed graphs, hypergraphs, time-varying networks) might be more appropriate would help guide methodological choices and avoid misapplication.
@@ Acknowledgements
We sincerely thank João Brázia, Violeta Calleja-Solanas, Gonzalo Contreras-Aso, Anna Gallo, Javier Garcia-Bernardo, Peter Gerbrands, Renaud Lambiotte, Francesco Paolo Nerini, Anxo Sánchez, and Mahdi Shafiee Kamalabad for their valuable feedback, with special thanks to Renaud Lambiotte for his extensive and insightful comments. We also gratefully acknowledge the discussions and panels at the FRIENDS Satellites of CCS 2024 and NetSci 2025, which inspired several parts of this review. F.D.-D. thanks funding from Program for Units of Excellences Maria de Maeztu (CEX2021-001164-M/10.13039/501100011033) and project APASOS (PID2021-122256NB-C22). M.A.G.-C. acknowledges support from the Comunidad de Madrid through the grants for the hiring of pre-doctoral research personnel in training (reference PIPF-2023/COM-29487). Both F.D.-D. and M.A.G.-C. thank grant PID2022-141802NB-I00 (BASIC) funded by MCIN/AEI/10.13039/501100011033 and by ‘ERDFA way of making Europe’. E.F. acknowledges funding from the Maria de Maeztu Units of Excellence Programme CEX2021-001195-M, funded by MICIU/AEI /10.13039/501100011033. I.F acknowledges support by Grant No. PRE2019-090279 (MCIN/AEI/10.13039/501100011033), Spanish Grant No. PID2021-128005NB-C22, funded by MCIN/AEI/ 10.13039/501100011033. A.S.T. acknowledges support by FCT – Fundação para a Ciência e Tecnologia – through the LASIGE Research Unit, ref. UID/408/2025.
@@ Appendix A Appendix
Kermarrec & Thraves (2014) [kermarrecSignedGraphEmbedding2014] Graph Drawing Euclidean embeddings aligning geometry with balance theory by placing positively connected nodes close together and negatively connected ones farther apart.
@@
Yuan et al. (2017) [yuanSNESignedNetwork2017] SNE Extension of the skip-gram objective to signed networks, enforcing proximity for friends and separation for enemies.
@@
Wang et al. (2017) [wangSignedNetworkEmbedding2017] SiNE Deep learning model for signed network embedding, enforcing structural balance with pairwise constraints. InzamamRahaman/SiNE
@@
Islam et al. (2018) [islamDistributedRepresentationsSigned2018] sign2vec Node2vec-style random walks adapted with edge signs and balance-aware context sampling. raihan2108/signet
@@
Javari et al. (2018) [javariStatisticalLinkLabel2018] STLGM Statistical latent graph model for sign prediction and embeddings in signed networks.
@@
Derr et al. (2018) [derrSignedGraphConvolutional2018] SGCN Graph convolutional network adapted for signed graphs, incorporating balance theory in aggregation. benedekrozemberczki/SGCN
@@
Schaub et al. (2019) [schaubMultiscaleDynamicalEmbeddings2019] Multiscale Dyn. Embeddings derived from diffusion dynamics across temporal and structural scales, capturing community and polarisation patterns. michaelschaub/DynamicalEmbeddings
@@
Huang et al. (2019) [huangSignedGraphAttention2019] SiGAT Signed graph attention networks learning embeddings by weighting neighbours with balance-aware attention. huangjunjie-cs/SiGAT
@@
Mara et al. (2020) [maraCSNEConditionalSigned2020] CSNE Conditional signed network embeddings using a maximum-entropy principle constrained by balance statistics. aida-ugent/CSNE
@@
Shen & Chung (2020) [shenDeepNetworkEmbedding2020] DNE-SBP Deep network embedding model for signed bipartite projections. shenxiaocam/Deep-network-embedding
@@
Song et al. (2021) [songHyperbolicNodeEmbedding2021] HSNE Hyperbolic embeddings on the Poincaré ball, well-suited for hierarchical and polarised signed structures.
@@
Yan et al. (2021) [yanMUSEMultifacetedAttention2021] MUSE Multifaceted attention signed GNN capturing different interaction types and their influence on embeddings.
@@
Huang et al. (2021) [huangSignedBipartiteGraph2021] SBGNN Signed bipartite graph neural network extending GNNs to heterogeneous signed settings. huangjunjie-cs/SBGNN
@@
Huang et al. (2022) [huangPOLEPolarizedEmbedding2022] POLE Random-walk–based embeddings preserving signed autocovariance, explicitly modelling multi-step polarisation. zexihuang/POLE
@@
He et al. (2022) [heSSSNETSemiSupervisedSigned2022] SSSNET Semi-supervised signed GNN for clustering and embeddings with balance constraints. SherylHYX/SSSNET_Signed_Clustering
@@
Pougué-Biyong et al. (2022) [pougue-biyongLearningStanceEmbeddings2022] SEM Stance embeddings combining signed node2vec walks with topical signals to capture ideological polarisation. lejohnyjohn/stance-embeddings
@@
Nakis et al. (2023) [nakisCharacterizingPolarizationSocial2023] SLDM-SLIM Latent distance embeddings mapping nodes to a simplex, representing polarisation through archetypal positions. Nicknakis/SLIM_RAA
@@
Nakis et al. (2023) [nakisHybridMembershipLatent2023] sHM-LDM Hybrid membership latent distance model allowing mixed affiliations and variable polarisation strength. Nicknakis/HM-LDM
@@
Chen et al. (2023) [chenBalancingAugmentationEdgeUtility2023] SiGAug Data augmentation strategy for robust signed GNN embeddings.
@@
Huang et al. (2023) [huangSDGNNLearningNode2023] SDGNN Signed dual GNN learning embeddings through dual message passing. huangjunjie-cs/SiGAT
Kim et al. (2023) [kimTrustworthinessDrivenGraphConvolutional2023] TrustSGCN Trustworthiness-driven signed GCN enhancing robustness and interpretability. kmj0792/TrustSGCN
@@
Sharma et al. (2023) [sharmaRepresentationLearningContinuousTime2023] SEMBA Continuous-time signed network embeddings via temporal point processes. claws-lab/semba
@@
Babul & Lambiotte (2024) [babulSHEEPSignedHamiltonian2024] SHEEP Spectral Hamiltonian embeddings based on a repelling Laplacian, emphasising separation of antagonistic groups. saynbabul/SHEEP
@@
Diaz-Diaz & Estrada (2025) [diazdiaz2025] communicability Communicability-based embeddings capturing signed geometric relationships. signed_graphs_via_communicability_geometry
@@
Guha et al. (2004) [guhaPropagationTrustDistrust2004] Structural balance propagation Sign inference from balance theory using products of signs along walks; negative influence propagates at most once.
Kunegis et al. (2009, 2010) [kunegisSlashdotZooMining2009, kunegisSpectralAnalysisSigned2010] Exponential and diffusion kernels Predicts signs using matrix powers of $A$ with penalization ($c_{k}=\beta^{k}$ or $1/k!$); variants include effective resistance and diffusion kernels.
Chiang et al. (2011) [chiangExploitingLongerCycles2011] Extended structural balance Incorporates higher-order cycles (up to $k$) into sign prediction, capturing longer balance patterns beyond triads.
Singh & Adhikari (2017) [singhMeasuringBalanceSigned2017] Walk-based balance Uses weighted sums of positive/negative walks to compute balance scores for edge signs.
Kirkley & Newman (2019) [kirkleyBalanceSignedNetworks2019] Balance-based inference Extends matrix-based formulations with factorial decays to estimate consistent sign assignments.
Leskovec et al. (2010) [leskovecPredictingPositiveNegative2010] Logistic regression Treats sign prediction as supervised classification; uses degree-based and triad-based structural features.
Chiang et al. (2011) [chiangExploitingLongerCycles2011] Logistic regression Improves Leskovec’s approach by incorporating cycle-based features (e.g., 4-cycles) for low-embeddedness edges.
Papaoikonomou et al. (2014) [papaoikonomouPredictingEdgeSigns2014] SVM with ego-graph motifs Uses frequent subgraph patterns (motifs) in local neighbourhoods as features for sign classification.
Derr et al. (2019) [derrBalanceSignedBipartite2019] Bipartite motif classifier Adapts motif-based features to bipartite signed networks (no triangles), using “caterpillars” and “butterflies.” DSE-MSU/signed-bipartite-networks
Forsati et al. (2015) [forsati2015pushtrust] Matrix factorization (MF) Learns latent vectors $U,V$ minimizing a sign-consistent loss; predicts sign by $\operatorname{sign}(u_{i}^{\top}v_{j})$.
Hsieh et al. (2012) [hsiehLowRankModeling2012] Low-rank matrix completion Recovers missing signs by completing the adjacency matrix under a low-rank constraint (using trace norm).
Ye et al. (2013) [ye_predicting_2013] Transfer matrix factorisation Jointly embeds source and target signed networks to transfer structural knowledge for improved sign prediction.
Ruiz-García et al. (2023) [ruiz-garciaTriadicInfluenceProxy2023] Neural imbalance predictor Neural network predicting edge signs from triadic influence; finds structural information more predictive than node attributes. miguel-rg/triadic-influence
Luo et al. (2021) [luoInterpretableSignedLink2021] SIHG Hyperbolic GNN with attention separating positive/negative neighbours, aligning with balance and status theories. Luoyadan/SIHG
Fang et al. (2023) [fangSignedSubgraphEncoding2023] SELO Deep encoder learning edge embeddings from subgraphs, capturing higher-order signed structures. fang98/SELO
Chen et al. (2023) [chenBalancingAugmentationEdgeUtility2023] SiGAug Graph augmentation framework filtering unreliable negatives to enhance robustness in GNN-based sign prediction.
@@
Doreian & Mrvar (1996)     [doreianPartitioningApproachStructural1996] Blockmodel Minimizes in-group disagreement and out-group agreement (Eq.     5). schochastics/signnet
Mrvar & Doreian (2009)     [mrvarPartitioningSignedTwoMode2009] Two-mode blockmodel Extension of blockmodels to bipartite signed networks. schochastics/signnet
Doreian & Mrvar (2009)     [doreianPartitioningSignedSocial2009] Relaxed blockmodel Generalization allowing more than two factions. schochastics/signnet
Gómez et al. (2009)     [gomezAnalysisCommunityStructure2009] Signed modularity Extension of modularity with separate null models for positive and negative links (Eq.     4.4.1).
Traag & Bruggeman (2009)     [traagCommunityDetectionNetworks2009] Spinglass Potts model Hamiltonian unifying frustration and modularity (eq.     6). igraph
Anchuri & Magdon-Ismail (2012)     [anchuriCommunitiesBalanceSigned2012] Double quality function Shows frustration minimisation and modularity maximisation are equivalent.
Amelio & Pizzuti (2013)     [amelioCommunityMiningSigned2013] SN-MOGA Multi-objective evolutionary algorithm optimising modularity and frustration.
Pougué-Biyong et al. (2024)     [pougue-biyongSignedLouvainLouvainSigned2024] SignedLouvain Modification of Louvain allowing exploration via negative links. lejohnyjohn/signed-louvain
Mendonça et al. (2015)     [mendoncaRelevanceNegativeLinks2015] Parallel ILS Iterative local search minimising frustration. CompNet/NetVotes
Aref et al. (2019, 2020)     [arefBalanceFrustrationSigned2019, arefModelingComputationalStudy2020, arefDetectingCoalitionsOptimally2020] Integer programming Exact optimization of frustration and coalition detection.
@@
Kunegis et al. (2010, 2014)     [kunegisSpectralAnalysisSigned2010, kunegisApplicationsStructuralBalance2014] Signed Laplacian Eigen-decomposition minimizing signed ratio cut (Eq.     7).
Chiang et al. (2012)     [chiangScalableClusteringSigned2012] Balance Ratio Cut Spectral relaxation handling multiple groups.
Cucuringu et al. (2019)     [cucuringuSPONGEGeneralizedEigenproblem2019] SPONGE Combines Laplacians of positive/negative subgraphs with parameters. alan-turing-institute/signet
Bonchi et al. (2019)     [bonchiDiscoveringPolarizedCommunities2019] EIGENSIGN Adjacency-based spectral method for polarised groups. egalimberti/polarized_communities
Zhong et al. (2020)     [zhongEfficientAlgorithmBased2020] SNBT Non-backtracking matrix eigen-decomposition for signed networks.
Fox et al. (2017, 2018)     [fox2017numerical, foxInvestigationSpectralClustering2018] Gremban expansion Embeds signed graphs into unsigned graphs for standard spectral clustering.
@@
Chen et al. (2014)     [chenOverlappingCommunityDetection2014] SPM Signed probabilistic mixture model, restrictive assumption of perfect balance.
Jiang (2015)     [jiangStochasticBlockModel2015] SBM variant Relaxes Chen’s model, allows positive/negative edges with different probabilities (Eq.     8).
Peixoto (2018)     [peixotoNonparametricWeightedStochastic2018] Weighted SBM Edge signs treated as covariates in hierarchical SBM. graph-tool.skewed.de
Aicher et al. (2015)     [aicherLearningLatentBlock2015] Generalised SBM Handles signed weights via exponential family distributions.
Yang et al. (2017)     [yangStochasticBlockmodelingVariational2017] Bayesian SBM Variational Bayesian inference for signed networks.
Zhao et al. (2017, 2018)     [zhaoStatisticalInferenceCommunity2017, zhaoNetworkModellingVariational2018] EM/Bayesian SBM EM-based and Bayesian approaches for signed networks.
Li et al. (2023)     [liSSBMSignedStochastic2023] SSBM Signed SBM with explicit sign probabilities.
Nakis et al. (2023)     [nakisHybridMembershipLatent2023] sHM-LDM Latent distance model embedding nodes in a simplex. Nicknakis/HM-LDM
Gallo et al. (2024)     [galloAssessingFrustrationRealworld2024] SSBM-BIC Maximum entropy SBM with BIC-based parameter selection.
@@
Hsieh et al. (2012)     [hsiehLowRankModeling2012] Matrix completion Low-rank modelling for signed adjacency matrices.
Esmailian et al. (2015)     [esmailianCommunityDetectionSigned2015] Infomap Extension of Infomap to signed networks. pouyaesm/signed-community-detection
Chu et al. (2016)     [chuFindingGangsWar2016] k-OCG Find $k$ cohesive subgraphs with high intra-graph cohesion and large inter-graph opposition.
Chang et al. (2017)     [changCommunityDetectionSigned2017] Hamming distance Breadth-first search, bit-flipping algorithm, and a belief propagation algorithm.
Su et al. (2017)     [suAlgorithmBasedPositive2017] SRWA Random walk strategies for signed networks.
Che et al. (2020)     [cheMemeticAlgorithmCommunity2020] MACD-SN Based on evolutionary theories, a memetic algorithm

### v2

@@ 8 Conclusions
Some fundamental quantities still lack accepted signed versions (e.g., assortativity). Others admit multiple possible generalisations: the Laplacian can be either opposing or repelling, node degree can be positive or negative depending on whether one considers signed, net, or total degree, and it is suitable for different contexts. In many cases, familiar properties break down when signs are introduced: shortest-path distances no longer satisfy metric axioms, and random-walk transition rates can become negative [383]. Even when a measure can be generalised, its interpretation often changes fundamentally. For instance, Katz centrality describes power relations and may become negative, and subgraph centrality encodes local balance rather than importance [90]. Developing principled frameworks for when to use which generalisation, and clarifying their distinct interpretations, would help standardise practice across the field.
@@
The term “polarisation” is used ambiguously in signed network research, encompassing distinct phenomena. Elite polarisation (in legislative or parliamentary contexts) differs from mass polarisation (in online platforms). Structural polarisation (the emergence of differentiated groups in relationship patterns) differs from affective polarisation (emotional or attitudinal hostility between groups) [199]. Moreover, signed networks enable distinguishing between mere separation, cohesive but disconnected groups, and explicit antagonism through negative ties. Understanding how these definitions relate and how empirical results differ when only endorsing actions are considered versus when negative interactions are also included is crucial [53], and what each reveals about underlying social or political processes remains an important open question.
@@
A foundational question remains often overlooked: what kinds of systems are best modelled as signed networks? While the framework is widely used, its application sometimes lacks clear justification. Not all oppositional or divergent relationships are best captured via edge signs. For example, there are serious concerns about treating correlation matrices as signed networks without prior processing [305]. Distinguishing between situations where signed networks provide genuine insight versus where other representations (multilayer networks, attributed graphs, hypergraphs, time-varying networks [5]) might be more appropriate would help guide methodological choices and avoid misapplication.
@@ Acknowledgements
We sincerely thank João Brázia, Violeta Calleja-Solanas, Gonzalo Contreras-Aso, Anna Gallo, Javier Garcia-Bernardo, Peter Gerbrands, Renaud Lambiotte, Francesco Paolo Nerini, Anxo Sánchez, and Mahdi Shafiee Kamalabad for their valuable feedback, with special thanks to Renaud Lambiotte for his extensive and insightful comments. We also gratefully acknowledge the discussions and panels at the FRIENDS Satellites of CCS 2024 and NetSci 2025, which inspired several parts of this review. We thank Nicole Samay for her careful work in refining the figures and ensuring a consistent visual style throughout the manuscript. F.D.-D. thanks funding from Program for Units of Excellences Maria de Maeztu (CEX2021-001164-M/10.13039/501100011033) and project APASOS (PID2021-122256NB-C22). M.A.G.-C. acknowledges support from the Comunidad de Madrid through the grants for the hiring of pre-doctoral research personnel in training (reference PIPF-2023/COM-29487). Both F.D.-D. and M.A.G.-C. thank grant PID2022-141802NB-I00 (BASIC) funded by MCIN/AEI/10.13039/501100011033 and by ‘ERDFA way of making Europe’. E.F. acknowledges funding from the Maria de Maeztu Units of Excellence Programme CEX2021-001195-M, funded by MICIU/AEI /10.13039/501100011033. I.F acknowledges support by Grant No. PRE2019-090279 (MCIN/AEI/10.13039/501100011033), Spanish Grant No. PID2021-128005NB-C22, funded by MCIN/AEI/ 10.13039/501100011033. A.S.T. acknowledges support by FCT – Fundação para a Ciência e Tecnologia – through the LASIGE Research Unit, ref. UID/408/2025.
@@ Appendix A Appendix
Kermarrec & Thraves (2014) [208] Graph Drawing Euclidean embeddings aligning geometry with balance theory by placing positively connected nodes close together and negatively connected ones farther apart.
@@
Yuan et al. (2017) [409] SNE Extension of the skip-gram objective to signed networks, enforcing proximity for friends and separation for enemies.
@@
Wang et al. (2017) [393] SiNE Deep learning model for signed network embedding, enforcing structural balance with pairwise constraints. InzamamRahaman/SiNE
@@
Islam et al. (2018) [198] sign2vec Node2vec-style random walks adapted with edge signs and balance-aware context sampling. raihan2108/signet
@@
Javari et al. (2018) [201] STLGM Statistical latent graph model for sign prediction and embeddings in signed networks.
@@
Derr et al. (2018) [86] SGCN Graph convolutional network adapted for signed graphs, incorporating balance theory in aggregation. benedekrozemberczki/SGCN
@@
Schaub et al. (2019) [343] Multiscale Dyn. Embeddings derived from diffusion dynamics across temporal and structural scales, capturing community and polarisation patterns. michaelschaub/DynamicalEmbeddings
@@
Huang et al. (2019) [190] SiGAT Signed graph attention networks learning embeddings by weighting neighbours with balance-aware attention. huangjunjie-cs/SiGAT
@@
Mara et al. (2020) [259] CSNE Conditional signed network embeddings using a maximum-entropy principle constrained by balance statistics. aida-ugent/CSNE
@@
Shen & Chung (2020) [352] DNE-SBP Deep network embedding model for signed bipartite projections. shenxiaocam/Deep-network-embedding
@@
Song et al. (2021) [360] HSNE Hyperbolic embeddings on the Poincaré ball, well-suited for hierarchical and polarised signed structures.
@@
Yan et al. (2021) [401] MUSE Multifaceted attention signed GNN capturing different interaction types and their influence on embeddings.
@@
Huang et al. (2021) [189] SBGNN Signed bipartite graph neural network extending GNNs to heterogeneous signed settings. huangjunjie-cs/SBGNN
@@
Huang et al. (2022) [192] POLE Random-walk–based embeddings preserving signed autocovariance, explicitly modelling multi-step polarisation. zexihuang/POLE
@@
He et al. (2022) [175] SSSNET Semi-supervised signed GNN for clustering and embeddings with balance constraints. SherylHYX/SSSNET_Signed_Clustering
@@
Pougué-Biyong et al. (2022) [314] SEM Stance embeddings combining signed node2vec walks with topical signals to capture ideological polarisation. lejohnyjohn/stance-embeddings
@@
Nakis et al. (2023) [284] SLDM-SLIM Latent distance embeddings mapping nodes to a simplex, representing polarisation through archetypal positions. Nicknakis/SLIM_RAA
@@
Nakis et al. (2023) [285] sHM-LDM Hybrid membership latent distance model allowing mixed affiliations and variable polarisation strength. Nicknakis/HM-LDM
@@
Chen et al. (2023) [65] SiGAug Data augmentation strategy for robust signed GNN embeddings.
@@
Huang et al. (2023) [191] SDGNN Signed dual GNN learning embeddings through dual message passing. huangjunjie-cs/SiGAT
Kim et al. (2023) [210] TrustSGCN Trustworthiness-driven signed GCN enhancing robustness and interpretability. kmj0792/TrustSGCN
@@
Sharma et al. (2023) [350] SEMBA Continuous-time signed network embeddings via temporal point processes. claws-lab/semba
@@
Babul & Lambiotte (2024) [28] SHEEP Spectral Hamiltonian embeddings based on a repelling Laplacian, emphasising separation of antagonistic groups. saynbabul/SHEEP
@@
Diaz-Diaz & Estrada (2025) [92] communicability Communicability-based embeddings capturing signed geometric relationships. signed_graphs_via_communicability_geometry
@@
Guha et al. (2004) [157] Structural balance propagation Sign inference from balance theory using products of signs along walks; negative influence propagates at most once.
Kunegis et al. (2009, 2010) [223, 224] Exponential and diffusion kernels Predicts signs using matrix powers of $A$ with penalization ($c_{k}=\beta^{k}$ or $1/k!$); variants include effective resistance and diffusion kernels.
Chiang et al. (2011) [68] Extended structural balance Incorporates higher-order cycles (up to $k$) into sign prediction, capturing longer balance patterns beyond triads.
Singh & Adhikari (2017) [356] Walk-based balance Uses weighted sums of positive/negative walks to compute balance scores for edge signs.
Kirkley & Newman (2019) [212] Balance-based inference Extends matrix-based formulations with factorial decays to estimate consistent sign assignments.
Leskovec et al. (2010) [231] Logistic regression Treats sign prediction as supervised classification; uses degree-based and triad-based structural features.
Chiang et al. (2011) [68] Logistic regression Improves Leskovec’s approach by incorporating cycle-based features (e.g., 4-cycles) for low-embeddedness edges.
Papaoikonomou et al. (2014) [303] SVM with ego-graph motifs Uses frequent subgraph patterns (motifs) in local neighbourhoods as features for sign classification.
Derr et al. (2019) [85] Bipartite motif classifier Adapts motif-based features to bipartite signed networks (no triangles), using “caterpillars” and “butterflies.” DSE-MSU/signed-bipartite-networks
Forsati et al. (2015) [123] Matrix factorization (MF) Learns latent vectors $U,V$ minimizing a sign-consistent loss; predicts sign by $\operatorname{sign}(u_{i}^{\top}v_{j})$.
Hsieh et al. (2012) [188] Low-rank matrix completion Recovers missing signs by completing the adjacency matrix under a low-rank constraint (using trace norm).
Ye et al. (2013) [408] Transfer matrix factorisation Jointly embeds source and target signed networks to transfer structural knowledge for improved sign prediction.
Ruiz-García et al. (2023) [332] Neural imbalance predictor Neural network predicting edge signs from triadic influence; finds structural information more predictive than node attributes. miguel-rg/triadic-influence
Luo et al. (2021) [251] SIHG Hyperbolic GNN with attention separating positive/negative neighbours, aligning with balance and status theories. Luoyadan/SIHG
Fang et al. (2023) [117] SELO Deep encoder learning edge embeddings from subgraphs, capturing higher-order signed structures. fang98/SELO
Chen et al. (2023) [65] SiGAug Graph augmentation framework filtering unreliable negatives to enhance robustness in GNN-based sign prediction.
@@
Doreian & Mrvar (1996)     [96] Blockmodel Minimizes in-group disagreement and out-group agreement (Eq.     5). schochastics/signnet
Mrvar & Doreian (2009)     [280] Two-mode blockmodel Extension of blockmodels to bipartite signed networks. schochastics/signnet
Doreian & Mrvar (2009)     [97] Relaxed blockmodel Generalization allowing more than two factions. schochastics/signnet
Gómez et al. (2009)     [147] Signed modularity Extension of modularity with separate null models for positive and negative links (Eq.     4.4.1).
Traag & Bruggeman (2009)     [386] Spinglass Potts model Hamiltonian unifying frustration and modularity (eq.     6). igraph
Anchuri & Magdon-Ismail (2012)     [13] Double quality function Shows frustration minimisation and modularity maximisation are equivalent.
Amelio & Pizzuti (2013)     [11] SN-MOGA Multi-objective evolutionary algorithm optimising modularity and frustration.
Pougué-Biyong et al. (2024)     [315] SignedLouvain Modification of Louvain allowing exploration via negative links. lejohnyjohn/signed-louvain
Mendonça et al. (2015)     [267] Parallel ILS Iterative local search minimising frustration. CompNet/NetVotes
Aref et al. (2019, 2020)     [22, 19, 20] Integer programming Exact optimization of frustration and coalition detection.
@@
Kunegis et al. (2010, 2014)     [224, 225] Signed Laplacian Eigen-decomposition minimizing signed ratio cut (Eq.     7).
Chiang et al. (2012)     [69] Balance Ratio Cut Spectral relaxation handling multiple groups.
Cucuringu et al. (2019)     [77] SPONGE Combines Laplacians of positive/negative subgraphs with parameters. alan-turing-institute/signet
Bonchi et al. (2019)     [46] EIGENSIGN Adjacency-based spectral method for polarised groups. egalimberti/polarized_communities
Zhong et al. (2020)     [419] SNBT Non-backtracking matrix eigen-decomposition for signed networks.
Fox et al. (2017, 2018)     [127, 128] Gremban expansion Embeds signed graphs into unsigned graphs for standard spectral clustering.
@@
Chen et al. (2014)     [66] SPM Signed probabilistic mixture model, restrictive assumption of perfect balance.
Jiang (2015)     [202] SBM variant Relaxes Chen’s model, allows positive/negative edges with different probabilities (Eq.     8).
Peixoto (2018)     [308] Weighted SBM Edge signs treated as covariates in hierarchical SBM. graph-tool.skewed.de
Aicher et al. (2015)     [3] Generalised SBM Handles signed weights via exponential family distributions.
Yang et al. (2017)     [403] Bayesian SBM Variational Bayesian inference for signed networks.
Zhao et al. (2017, 2018)     [418, 417] EM/Bayesian SBM EM-based and Bayesian approaches for signed networks.
Li et al. (2023)     [238] SSBM Signed SBM with explicit sign probabilities.
Nakis et al. (2023)     [285] sHM-LDM Latent distance model embedding nodes in a simplex. Nicknakis/HM-LDM
Gallo et al. (2024)     [139] SSBM-BIC Maximum entropy SBM with BIC-based parameter selection.
@@
Hsieh et al. (2012)     [188] Matrix completion Low-rank modelling for signed adjacency matrices.
Esmailian et al. (2015)     [105] Infomap Extension of Infomap to signed networks. pouyaesm/signed-community-detection
Chu et al. (2016)     [71] k-OCG Find $k$ cohesive subgraphs with high intra-graph cohesion and large inter-graph opposition.
Chang et al. (2017)     [62] Hamming distance Breadth-first search, bit-flipping algorithm, and a belief propagation algorithm.
Su et al. (2017)     [368] SRWA Random walk strategies for signed networks.
Che et al. (2020)     [64] MACD-SN Based on evolutionary theories, a memetic algorithm